OpenStudy (mazehaq):

Find the exact value of 2sec(π/4) + 4cot(π/3). Do I find cos(π/4) & tan(π/3)?

OpenStudy (anonymous):

yes...

OpenStudy (mazehaq):

cos(π/4)= sqrt2/2 & tan(π/3)= sqrt3, right?

OpenStudy (anonymous):

yes... so just take the reciprocals and replace them in your original problem.

OpenStudy (mazehaq):

Do I ignore sec & cot, sorry, I just learned this today.

OpenStudy (anonymous):

no... \(\large cos\frac{\pi}{4}=\frac{\sqrt2}{2} \) so this means: \(\large sec\frac{\pi}{4}=\frac{2}{\sqrt2} \)

OpenStudy (mazehaq):

Right, I got that.

OpenStudy (mazehaq):

So far.

OpenStudy (anonymous):

and since \(\large tan\frac{\pi}{3}=\sqrt3 \) then \(\large cot\frac{\pi}{3}=\frac{1}{\sqrt3} \) plug these values into the original problem

OpenStudy (anonymous):

\(\large 2sec\frac{\pi}{4}+4cot\frac{\pi}{3}=2(\frac{2}{\sqrt2})+4(\frac{1}{\sqrt3}) \)

OpenStudy (anonymous):

first one is (1/ (cos (pie/4)))*2 = 2 *under root 2 second one (1/tan(pie/3))*4 =4 under root 3 5.14 is the answer :D

OpenStudy (mazehaq):

Thanks you guys, but ByteMe, wouldn't I rationalize & then solve?

OpenStudy (anonymous):

also divide by 3